An isosceles right triangle has a perimeter of 14 cm. Find the length of a leg and simplify your answer. Leave answers with exact roots, do not use calculator!
Answer by lwsshak3(11628)

An isosceles right triangle has a perimeter of 14 cm. Find the length of a leg and simplify your answer. Leave answers with exact roots, do not use calculator!
***
let x=length of each leg
by the pythagorean theorem:
hypotenuse=√x^2+x^2=√2x^2=√2x
x+x+√2x=14
2x+√2x=14
x(2+√2)=14
x=14/(2+√2)
length of each leg=14/(2+√2)
check:
x+x+√2x=2x+√2x=x(2+√2)=14/(2+√2)*(2+√2)=14